What is the volume of aluminum ball at 10⁰C it it's volume at 160⁰C is 300cm³?


1
Expert's answer
2021-05-04T12:07:30-0400

"\frac{L -L_0}{L_0} = α_VΔT \\

L = L_0(1+α_VΔT ) \\

L = L_0(1 + 3α_LΔT ) \\

L = 300(1 + 3 \times 23 \times 10^{-6}(10 -160)) = 297 \;m^3"